# Neural Scaling Laws for Deep Regression

### v1

@@ I Introduction
Figure 1: Neural scaling laws in a parameter estimation model. Panels (a)–(c) show the geometric mean of mean squared error (MSE) test loss in the prediction of three magnetic parameters ($\theta$, $J$, $D$), respectively, as a function of training dataset size. Each marker corresponds to a different architecture: fully connected networks (FCNs), residual network (ResNet-18), and vision transformer (ViT). The two FCNs, indicated by red and yellow squares, have different numbers of neurons per hidden layer ($n_{n}=4$ and $n_{n}=512$, respectively) with the same number of hidden layers ($n_{l}=3$). In each panel, dashed lines represent power-law fits of the form $\epsilon\sim N_{D}^{-\alpha_{D}}$, where $\epsilon$ is the loss, $N_{D}$ is the training dataset size, and $\alpha_{D}$ is the scaling exponent. All data are plotted on a log-log scale, with dataset sizes ranging from 224 to 114688. Data points for the two largest model sizes in panel (a) are omitted.
@@
Deep learning has established a new data-driven programming paradigm, where data determines the algorithms used to solve specific problems. This approach often provides substantial advantages over traditional logic-driven programming, particularly in addressing intricate scientific and real-world problems, where the required logic of algorithms is unknown or too complex. Generally, the performance of deep learning models depends on the quantity and quality of the data used to train deep neural networks. Researchers have identified various techniques for enhancing model performance, either by utilizing various types of regularization and batch normalization from the model side, or by pre-processing a given dataset [1] or through data augmentation [2, 3] from the data side. The latter potential can be systematically examined using “generalization error scaling laws,” which describe the power-law relationship between generalization errors and dataset size, model size and the computing time used for the training.
@@
Early theoretical studies sought to derive generalization error scaling laws by solving statistical models [4, 5], while recent research has focused on empirical investigations in the context of classification models [6, 7, 8, 9, 10, 11, 12, 13, 14, 12, 15, 16, 17] and language models [18, 16, 19, 16, 15, 18, 20, 16, 17]. In particular, the exploration of the scaling laws in large language models found that the sizes of datasets and models are crucial for model performance, while other features such as network structure have secondary impacts [18, 17]. These insights have significantly contributed to the recent success of large language models. In contrast to substantial advancements in classification and language models, the scaling laws for deep regression models remain largely understudied, despite recent emerging research [21]. Gaining a deeper understanding of these laws could provide valuable insights for developing deep regression models with predictable performance. Such insights are particularly crucial in scientific disciplines, where computational and data resources are often limited and optimizing resource usage is important.
@@
In this work, we empirically investigate the scaling laws of generalization error in deep regression, using a parameter estimation model that predicts magnetic Hamiltonian parameters from input magnetic domain images [Fig. 2]. By training networks on datasets of varying sizes or on multiple networks with a fixed dataset, we examine how the mean squared error (MSE) test loss in the model’s predictions scales with the training dataset size. We find that the test loss follows a power-law scaling behavior, expressed as $\epsilon\sim N_{D}^{-\alpha_{D}}$, where $\epsilon$ is the MSE test loss, $N_{D}$ is the dataset size, and $\alpha_{D}$ is the corresponding scaling exponent [Fig. 1]. Notably, the scaling behavior often spans more than three orders of magnitude in dataset size. Furthermore, the scaling exponents $\alpha_{D}$ range from approximately 1 to 2—much larger than those typically observed in classification or language models, which are often below 1 [6, 7, 8, 9, 11, 10, 12, 13, 14, 12, 15, 16, 18, 19, 20], with a rare case slightly above 1 [17]. This suggests that the model’s performance improves more rapidly with increased data or model complexity than typically expected. Additionally, increasing model complexity significantly reduces the MSE test loss, showing a similar power-law scaling $\epsilon\sim N_{M}^{-\alpha_{M}}$ with the model size $N_{M}$ and the scaling exponent $\alpha_{M}$ [Fig. 6]. These findings demonstrate that deep regression models can achieve substantial improvements in accuracy as data and model sizes grow.
@@ II.1 Neural network realizations
Figure 3: Statistics of MSE test loss across different realizations of FCNs. (a)–(d) Distributions of test loss across multiple network realizations for the regression of $J$, with bars indicating the number of networks $N_{\mathrm{bin}}$ within each loss range. Each panel corresponds to a specific combination of the number of neurons per hidden layer $n_{n}$ and dataset size $N_{D}$, with a fixed number of hidden layers $n_{l}=3$: panels (a) and (c) show results for $n_{n}=4$ and $N_{D}=3584$; panels (b) and (d) are for $n_{n}=16$ and $N_{D}=28672$. Panels (a) and (b) display distributions over 20 network realizations, whereas panels (c) and (d) show results over 100 realizations. In each panel, black, blue, and red lines denote the arithmetic mean, geometric mean, and median of each distribution, respectively, with shaded regions representing the standard errors of the arithmetic and geometric means and the median absolute deviation, each depicted in their corresponding colors. (e) Evolution of the geometric mean with increasing number of network realizations. Results for four model sizes ($n_{n}=4,16,64,256$) are shown in red, orange, green, and blue. For each size, data are presented across three testing dataset sizes $N_{D}=1792$, 7168, and 28672, marked by square, pentagon, and hexagon symbols, respectively. Error bars indicate the standard deviation of the geometric means across bootstrap samples. The bootstrap procedure involves generating 50 random subsets from the data of 100 networks, computing the geometric mean for each subset, and plotting the average of these means.
@@
Our atomistic spin simulations of twisted bilayer CrI3 [22, 23, 24, 25, 26] generated 162782 pairs of simulated magnetic domain images. These magnetic domain images exhibit regular yet complex variation patterns in response to changes in the underlying spin model parameters [27, 28, 22, 23, 24, 26], as illustrated in Fig. 2(a). The parameter estimation model employed in this study takes local out-of-plane magnetization, represented as two-dimensional arrays, as input, and predicts the underlying magnetic Hamiltonian parameters $(\theta,J,D)$ as output [Fig. 2(b)]. Using various neural network architectures—including fully-connected networks (FCNs), ResNet-18 [29], and a vision transformer (ViT) [30]—we investigate how the MSE test loss in parameter prediction scales with the training dataset size. We use MSE test loss as our primary metric to measure the network’s performance throughout this work. For FCNs, we examine a range of model sizes by varying either the number of neurons per hidden layer (with a fixed number of hidden layers) or the number of hidden layers (with a fixed number of neurons in each layer) to investigate how the model size influences the scaling behavior of the test loss. The training datasets are randomly sampled, and the values of the test loss are averaged over multiple network realizations. A consistent test dataset is used throughout this study. Further details on the atomistic spin simulations and deep learning methods can be found in the Methods section.
@@
Figures 3(a)–(d) illustrate the test loss distributions among various network realizations within the same network architecture and training dataset size. The test loss distributions are not uniform due to the intrinsic stochasticity of random sampling and the inherent randomness in the network optimization process [31]. The non-uniformity of the distribution is particularly significant in smaller models where outliers with relatively poor performance incidentally appear [Figs. 3(a) and (c)]. The presence of outliers biases the arithmetic mean, rendering it an unreliable measure of average performance, especially in smaller models. In contrast, the geometric mean exhibits greater robustness to outliers and provides a more accurate representation of the test loss distribution. Accordingly, we adopt the geometric mean to evaluate the average performance across different network realizations. Using the geometric mean provides a consistent framework for assessing model performance across various network sizes.
@@
Figure 3(e) shows the geometric mean of the test loss as a function of the number of trained networks used in averaging. The geometric mean plateaus once a sufficiently large number of network realizations is included. Increasing the number of realizations also decreases the variance, indicating that reliable estimates of the average performance can be achieved with a large enough sample size. In our analysis, we utilize twenty network realizations to estimate the geometric mean of the test loss, providing a robust measure across all considered architectures. Notably, larger models tend to have smaller variances, and increasing the dataset size further reduces the variance. In other words, different network realizations in the large model and dataset regimes demonstrate consistent results. This uniformity suggests that the derived scaling laws are more reliable and possess strong predictive power for real-world applications, where constructing only a limited number of models is typical, unlike the extensive stochastic averaging employed in this study.
@@
We found that accurately capturing the scaling behavior of the test loss requires careful tuning of the learning rate during training. For large models, using a fixed learning rate results in substantial fluctuations in test loss as a function of dataset size, rather than the expected power-law relationships. In contrast, gradually decreasing the learning rate enables optimization of model performance across different dataset sizes, thereby revealing well-established scaling behavior. This adaptive learning rate strategy was applied consistently across all network realizations; detailed method is presented in the Methods section. Furthermore, we tested standard regularization schemes such as L2 regularization and dropout; however, these techniques did not enhance the consistency of the observed scaling behaviors. Detailed analyses are provided in the Supplementary Information.
@@ II.2 Scaling laws with the data set size
Figure 4: Geometric mean of MSE test loss as a function of training dataset size. Panels (a)–(c) show results for the regression of $\theta$, $J$, and $D$, respectively. Each marker represents FCNs with varying $n_{n}$ from 4 to 512 with a fixed value of $n_{l}=3$. Panel (d) presents the result for the regression of $J$, with each marker indicating different values of $n_{l}$ and a fixed value of $n_{n}=16$. In each panel, dashed lines are power-law fits of the form $\epsilon\sim N_{D}^{-\alpha_{D}}$. In panel (d), the fitted values of $\alpha_{D}$ are included in the legend. All data are plotted on a log-log scale, with dataset sizes ranging from 224 to 114688. The green square indicates the result reported in Ref. 25.
@@
Figure 4 illustrates the variation of the test loss in predicting magnetic Hamiltonian parameters as a function of training dataset size across different FCN architectures. The test loss displays a monotonically decreasing trend with increasing dataset size for all regressed parameters and network architectures. These trends are well captured by a power-law scaling formula $\epsilon\sim N_{D}^{-\alpha_{D}}$, as indicated by the dashed lines. Notably, this power-law scaling is most evident in the small to intermediate dataset regimes (up to 28,672, as shown in the plots), spanning several orders of magnitude in both test loss and dataset size. Beyond this regime, the scaling breaks down. For $\theta$, the test loss exhibits substantial fluctuations in this regime [Fig. 4(a)], likely due to the numerical precision limit of approximately $1.9\times 10^{-7}$ imposed by the 32-bit floating-point calculations used in our computations, thus values below this are considered unreliable. In contrast, for parameters $J$ and $D$, the loss displays slowing down from the trend [Figs. 4(b) and (c)]. It is worth noting that the FCN with approximately $3.2\times 10^{5}$ parameters achieves a test loss comparable to that reported in Ref. 25 (which used an FCN with approximately $2.1\times 10^{7}$ parameters) for the regression of $J$ and $D$, while showing a substantial deviation for $\theta$.
@@
Figure 4(d) illustrates the impact of varying the number of hidden layers, considering FCN architectures with 1 to 5 layers. Across all configurations, the test loss demonstrates clear power-law scaling with respect to the training dataset size. However, the power-law exponents differ by network depth: shallower networks ($n_{l}=1,2$) exhibit smaller exponents, indicating a reduced capacity to improve performance as dataset size increases. In contrast, deeper networks ($n_{l}=3,4,5$) show nearly saturated exponents with only marginal gains from added layers, suggesting diminishing returns in scaling benefits beyond a certain depth. This behavior aligns with findings in the literature, where the scaling exponent increases with model complexity but tends to saturate as networks become deeper, reflecting intrinsic limitations in optimization and model expressiveness for practical dataset sizes [11, 17].
@@
Figure 5 shows the power-law scaling exponents $\alpha_{D}$, derived from the test loss data in Fig. 4 by fitting to $\epsilon\sim N_{D}^{-\alpha_{D}}$. The values of $\alpha_{D}$ vary considerably with model size: ranging from 1.4 to 2.3 for $\theta$, 0.8 to 1.8 for $J$, and 1.2 to 1.8 for $D$. Notably, the exponents tend to increase with model size. The logarithmic fit $\alpha_{D}=a_{D}\ln(N_{M})+b_{D}$ matches the data well for $J$, while it is less pronounced for $D$ and $\theta$. The fitted parameters are: $a_{D}=0.124\pm 0.048,b_{D}=0.30\pm 0.65$ for $\theta$; $a_{D}=0.184\pm 0.022,b_{D}=-1.17\pm 0.30$ for $J$; and $a_{D}=0.112\pm 0.042,b_{D}=0.06\pm 0.56$ for $D$. Large error bars are observed primarily for $\theta$, making conclusions for this parameter less definitive. Additionally, the exponents for $\theta$ and $D$ decline at the largest dataset sizes, deviating from the overall increasing trend. This likely results from a strong dependence of the test loss on the training scheme, suggesting potential for further network optimization. Therefore, more refined analysis may be needed to accurately characterize behavior in this large dataset regime. For comparison, the exponents for ResNet-18 and ViT, derived from the test loss data in Fig. 1, are also displayed in Fig. 5. These do not follow the same logarithmic pattern, despite involving the same dataset and task.
@@
We observed that the average of MSE test loss, computed using either the arithmetic mean or median, exhibits scaling behaviors similar to the geometric mean: in both cases, the test loss decreases monotonically with increasing dataset size, with more pronounced reductions for larger models. Their scaling exponents show quantitative agreement with those from the geometric mean. This consistency can be attributed to the convergence of the geometric mean, arithmetic mean, and median for sufficiently large models [Figs. 3(b) and (d)]. Consequently, these measures can be used interchangeably for large models. However, adopting the geometric mean provides a more robust and consistent framework for evaluating model performance across all network sizes. Further details are available in the Supplementary Information.
@@
In summary, we find that, for a given network architecture, the test loss exhibits well-defined power-law scaling with dataset size $N_{D}$, and increasing model complexity generally leads to higher scaling exponents. However, the relationship between the exponent and model complexity depends on detailed network architecture features beyond just the number of parameters, contrasting with previous findings [18].
@@
Figure 5: Power-law scaling exponents $\alpha_{D}$ as a function of FCN model size. Each marker displays the fitted parameter of $\alpha_{D}$ from the MSE test loss data in Fig. 4(a)–(c) for $\theta$, $J$, or $D$, respectively. The FCN model size $N_{M}$ ranges from approximately $8\times 10^{4}$ to $10^{7}$ parameters, corresponding to FCNs with $n_{n}$ from 4 to 512 and $n_{l}=3$. Dashed lines represent logarithmic fits of the form $\alpha_{D}=a_{D}\ln(N_{M})+b_{D}$, where $a_{D}$ and $b_{D}$ are fitting parameters. The data points for the largest value of $N_{M}$ are omitted from the fit. Circle and diamond markers indicate the $\alpha_{D}$ values for ResNet-18 and ViT, respectively.

### v2

@@ I Introduction
Figure 1: Neural scaling laws in a parameter estimation model. Panels (a)–(c) show the geometric mean of the mean squared error (MSE) test loss in the prediction of three magnetic parameters ($\theta$, $J$, $D$), respectively, as a function of training dataset size. Each marker corresponds to a different architecture: fully connected networks (FCNs), residual network (ResNet-18), and vision transformer (ViT). The two FCNs, indicated by red and yellow squares, have different numbers of neurons per hidden layer ($n_{n}=4$ and $n_{n}=512$, respectively) with the same number of hidden layers ($n_{l}=3$). In each panel, dashed lines represent power-law fits of the form $\epsilon\sim N_{D}^{-\alpha_{D}}$, where $\epsilon$ is the loss, $N_{D}$ is the training dataset size, and $\alpha_{D}$ is the scaling exponent. All data are plotted on a log-log scale, with dataset sizes ranging from 224 to 114688.
@@
Deep learning has established a new data-driven programming paradigm, where data determines the algorithms used to solve specific problems. This approach often provides substantial advantages over traditional logic-driven programming, particularly in addressing intricate scientific and real-world problems, where the required logic of algorithms is unknown or too complex. Generally, the performance of deep learning models depends on the quantity and quality of the data used to train deep neural networks. Researchers have identified various techniques for enhancing model performance, either by utilizing various types of regularization and batch normalization from the model side, or by pre-processing a given dataset [8686063] or through data augmentation [perez2017effectiveness, Steiner2022HowViT] from the data side. The latter potential can be systematically examined using “generalization error scaling laws,” which describe the power-law relationship between generalization errors and dataset size, model size and the computing time used for the training.
@@
Several early [AMARI1993161, 6796972] as well as more recent [maloney2022solvable, Defilippis2023DimensionFree, Paquette2024phases, bahri2024explaining] theoretical studies sought to derive generalization error scaling laws by solving statistical models, while recent research has focused on empirical investigations in the context of classification models [BAILLY2022106504, hestness2017deeplearningscalingpredictable, rosenfeld2020constructive, henighan2020scaling, mühlenstädt2024dataneed2predicting, meir2020powerlaw, Figueroa2012, cho2016dataneededtrainmedical, 8237359, Figueroa2012, JMLR:v23:20-1111, NEURIPS2022_8c22e5e9] and language models [kaplan2020scalinglawsneurallanguage, NEURIPS2022_8c22e5e9, hoffmann2022trainingcomputeoptimallargelanguage, NEURIPS2022_8c22e5e9, JMLR:v23:20-1111, kaplan2020scalinglawsneurallanguage, hutter2021learningcurvetheory, NEURIPS2022_8c22e5e9]. In particular, the exploration of the scaling laws in large language models found that the sizes of datasets and models are crucial for model performance, while other features such as network structure have secondary impacts [kaplan2020scalinglawsneurallanguage, bahri2024explaining]. These insights have significantly contributed to the recent success of large language models. In contrast to substantial advancements in classification and language models, the scaling laws for deep regression models remain largely understudied, despite recent emerging research [lin2024scalinglawslinearregression]. Gaining a deeper understanding of these laws could provide valuable insights for developing deep regression models with predictable performance. Such insights are particularly crucial in scientific disciplines where computational and data resources are often limited and optimizing resource usage is important.
@@
In this work, we empirically investigate the scaling laws of generalization error in deep regression, using a parameter estimation model that predicts magnetic Hamiltonian parameters from input magnetic domain images [Fig. 2]. By training networks on datasets of varying sizes or on multiple networks with a fixed dataset, we examine how the mean squared error (MSE) test loss in the model’s predictions scales with the training dataset size. We find that the test loss follows a power-law scaling behavior, expressed as $\epsilon\sim N_{D}^{-\alpha_{D}}$, where $\epsilon$ is the MSE test loss, $N_{D}$ is the dataset size, and $\alpha_{D}$ is the corresponding scaling exponent [Fig. 1]. Notably, the scaling behavior often spans more than three orders of magnitude in dataset size. Furthermore, the scaling exponents $\alpha_{D}$ range from approximately 1 to 2—much larger than those typically observed in classification or language models, which are often below 1 [BAILLY2022106504, hestness2017deeplearningscalingpredictable, rosenfeld2020constructive, henighan2020scaling, meir2020powerlaw, mühlenstädt2024dataneed2predicting, Figueroa2012, cho2016dataneededtrainmedical, 8237359, Figueroa2012, JMLR:v23:20-1111, NEURIPS2022_8c22e5e9, kaplan2020scalinglawsneurallanguage, hoffmann2022trainingcomputeoptimallargelanguage, hutter2021learningcurvetheory], with a rare case slightly above 1 [bahri2024explaining] and a recent study [maloney2022solvable] with exponents of up to 2. This suggests that the model’s performance improves more rapidly with increased data or model complexity than typically expected. Additionally, increasing model complexity significantly reduces the MSE test loss, showing a similar power-law scaling $\epsilon\sim N_{M}^{-\alpha_{M}}$ with the model size $N_{M}$ and the scaling exponent $\alpha_{M}$ [Fig. 6]. These findings demonstrate that deep regression models can achieve substantial improvements in accuracy as data and model sizes grow.
@@ II.1 Neural network realizations
Figure 3: Statistics of MSE test loss across different realizations of FCNs. (a)–(d) Distributions of test loss across multiple network realizations $N_{r}$ for the regression of $J$, with bars indicating the number of networks $N_{\mathrm{bin}}$ within each loss range. Each panel corresponds to a specific combination of the number of neurons per hidden layer $n_{n}$ and dataset size $N_{D}$, with a fixed number of hidden layers $n_{l}=3$: panels (a) and (c) show results for $n_{n}=4$ and $N_{D}=3584$; panels (b) and (d) are for $n_{n}=16$ and $N_{D}=28672$. Panels (a) and (b) display distributions over $N_{r}=20$ network realizations, whereas panels (c) and (d) show results with $N_{r}=100$ realizations. In each panel, black, blue, and red lines denote the arithmetic mean, geometric mean, and median of each distribution, respectively, with shaded regions representing the standard errors of the arithmetic and geometric means and the median absolute deviation, each depicted in their corresponding colors. (e) Evolution of the geometric mean with increasing number of network realizations. Results for four model sizes ($n_{n}=4,16,64,256$) are shown in red, orange, green, and blue. For each size, data are presented across three testing dataset sizes $N_{D}=1792$, 7168, and 28672, marked by square, pentagon, and hexagon symbols, respectively. Error bars indicate the standard deviation of the geometric means across bootstrap samples. The bootstrap procedure involves generating 50 random subsets from the data of 100 networks, computing the geometric mean for each subset, and plotting the average of these means.
@@
Our atomistic spin simulations of twisted bilayer CrI3 [Kim2023, PhysRevB.108.L100401, Kim2024, Lee_2024, Kim2025] generated 162782 pairs of simulated magnetic domain images. These magnetic domain images exhibit regular yet complex variation patterns in response to changes in the underlying spin model parameters [PhysRevResearch.3.013027, Zheng2022, Kim2023, PhysRevB.108.L100401, Kim2024, Kim2025], as illustrated in Fig. 2(a). The parameter estimation model employed in this study takes local out-of-plane magnetization, represented as two-dimensional arrays, as input, and predicts the underlying magnetic Hamiltonian parameters $(\theta,J,D)$ as output [Fig. 2(b)]. Using various neural network architectures—including fully-connected networks (FCNs), ResNet-18 [He2016-ResNet], and a vision transformer (ViT) [dosovitskiy2021image]—we investigate how the MSE test loss in parameter prediction scales with the training dataset size. We use MSE test loss as our primary metric to measure the network’s performance throughout this work. For FCNs, we examine a range of model sizes by varying either the number of neurons per hidden layer (with a fixed number of hidden layers) or the number of hidden layers (with a fixed number of neurons in each layer) to investigate how the model size influences the scaling behavior of the test loss. The training datasets are randomly sampled, and the values of the test loss are averaged over multiple network realizations. Each realization is defined as an independently trained instance of the network, starting with a random initialization for its weights and biases. A consistent test dataset is used throughout this study. Further details on the atomistic spin simulations and deep learning methods can be found in the Methods section.
@@
Figures 3(a)–(d) illustrate the test loss distributions among various network realizations within the same network architecture and training dataset size. The test loss distributions are not uniform due to the intrinsic stochasticity of random sampling and the inherent randomness in the network optimization process [Mitchell1997]. The non-uniformity of the distribution is particularly significant in smaller models where outliers with relatively poor performance incidentally appear [Figs. 3(a) and (c)]. The presence of outliers biases the arithmetic mean, rendering it an unreliable measure of average performance, especially in smaller models. In contrast, the geometric mean exhibits greater robustness to outliers and provides a more accurate representation of the test loss distribution. Accordingly, we adopt the geometric mean to evaluate the average performance across different network realizations. Using the geometric mean provides a consistent framework for assessing model performance across various network sizes.
@@
Figure 3(e) shows the geometric mean of the test loss as a function of the number of trained networks used in averaging. The geometric mean for $n_{n}=\{16,64,256\}$ is relatively constant across the number of realizations. In contrast, the geometric mean for $n_{n}=4$ shows substantial fluctuations at small realization counts while plateaus once a sufficiently large number of network realizations is included. Increasing the number of realizations also decreases the variance, indicating that reliable estimates of the average performance can be achieved with a large enough sample size. In our analysis, we utilize twenty network realizations to estimate the geometric mean of the test loss, providing a robust measure across all considered architectures. Notably, larger models tend to have smaller variances, and increasing the dataset size further reduces the variance. In other words, different network realizations in the large model and dataset regimes demonstrate consistent results. This uniformity suggests that the derived scaling laws are more reliable and possess strong predictive power for real-world applications, where constructing only a limited number of models is typical, unlike the extensive stochastic averaging employed in this study.
@@
We found that accurately capturing the scaling behavior of the test loss requires careful tuning of the learning rate during training. For smaller models, a fixed learning rate ensures consistent power-law scaling up to the largest dataset size used. However, for large models, using a fixed learning rate results in substantial fluctuations in test loss as a function of dataset size, rather than the expected power-law relationships across a large dataset size regime. In contrast, gradually decreasing the learning rate enables optimization of model performance across different dataset sizes, thereby revealing well-established scaling behavior. A key difference between the fixed and adjustable learning rate schemes lies in the scaling exponent values and the applicable scaling regimes. To examine the power-law scaling laws across the largest dataset size regime, we employ an adjustable learning rate scheme throughout this work; the detailed methodology is presented in the Methods section. Furthermore, we tested standard regularization techniques such as L2 regularization and dropout; however, these approaches did not enhance the consistency of the observed scaling behaviors. Detailed analyses are provided in the Supplementary Information.
@@ II.2 Scaling laws with the data set size
Figure 4: Geometric mean of MSE test loss as a function of training dataset size. Panels (a)–(c) show results for the regression of $\theta$, $J$, and $D$, respectively. Each marker represents FCNs with varying $n_{n}$ from 4 to 512 with a fixed value of $n_{l}=3$. Panel (d) presents the result for the regression of $J$, with each marker indicating different values of $n_{l}$ and a fixed value of $n_{n}=16$. In each panel, dashed lines are power-law fits of the form $\epsilon\sim N_{D}^{-\alpha_{D}}$. In panel (d), the fitted values of $\alpha_{D}$ are included in the legend. All data are plotted on a log-log scale, with dataset sizes ranging from 224 to 114688. The green square in panels (a)–(c) indicates the result reported in Ref. Lee_2024.
@@
Figure 4 illustrates the variation of the test loss in predicting magnetic Hamiltonian parameters as a function of training dataset sizes across different FCN architectures. The test loss displays a monotonically decreasing trend with increasing dataset size for all regressed parameters and network architectures. These trends are well captured by a power-law scaling formula $\epsilon\sim N_{D}^{-\alpha_{D}}$, as indicated by the dashed lines. Notably, this power-law scaling is most evident in the small to intermediate dataset regimes (up to 28,672, as shown in the plots), spanning several orders of magnitude in both test loss and dataset size. Beyond this regime, the scaling breaks down. For $\theta$, the test loss exhibits substantial fluctuations in this regime [Fig. 4(a)], likely due to the numerical precision limit of approximately $1.9\times 10^{-7}$ imposed by the 32-bit floating-point calculations used in our computations of regressed parameters and corresponding MSE test loss, thus values below this are considered unreliable. In contrast, for parameters $J$ and $D$, the loss displays slowing down from the trend [Figs. 4(b) and (c)]. It is worth noting that the FCN with approximately $3.2\times 10^{5}$ parameters achieves a test loss comparable to that reported in Ref. Lee_2024 (which used an FCN with approximately $2.1\times 10^{7}$ parameters) for the regression of $J$ and $D$, while showing a substantial deviation for $\theta$.
@@
Figure 4(d) illustrates the impact of varying the number of hidden layers, considering FCN architectures with 1 to 5 layers. Across all configurations, the test loss demonstrates clear power-law scaling with respect to the training dataset size. However, the power-law exponents differ by network depth: shallower networks ($n_{l}=1,2$) exhibit smaller exponents, indicating a reduced capacity to improve performance as dataset size increases. In contrast, deeper networks ($n_{l}=3,4,5$) show nearly saturated exponents with only marginal gains from added layers, suggesting diminishing returns in scaling benefits beyond a certain depth. This behavior aligns with findings in the literature, where the scaling exponent increases with model complexity but tends to saturate as networks become deeper, reflecting intrinsic limitations in optimization and model expressiveness for practical dataset sizes [meir2020powerlaw, bahri2024explaining].
@@
Figure 5 shows the power-law scaling exponents $\alpha_{D}$, derived from the test loss data in Fig. 4 by fitting to $\epsilon\sim N_{D}^{-\alpha_{D}}$. The values of $\alpha_{D}$ vary considerably with model size: ranging from 1.4 to 2.3 for $\theta$, 0.8 to 1.8 for $J$, and 1.2 to 1.8 for $D$. Notably, the exponents tend to increase with model size. The logarithmic fit for the dataset size exponent, $\alpha_{D}=a_{D}\log_{10}(N_{M}/10^{6})+b_{D}$, matches the data well for $J$, while it is less pronounced for $D$ and $\theta$; we emphasize that here the subscript $D$ is associated with the dataset size not the magnetic parameter $D$. The fitted parameters are: $a_{D}=0.28\pm 0.11,b_{D}=2.00\pm 0.07$ for $\theta$; $a_{D}=0.42\pm 0.05,b_{D}=1.38\pm 0.03$ for $J$; and $a_{D}=0.26\pm 0.10,b_{D}=1.61\pm 0.06$ for $D$. Large error bars are observed primarily for $\theta$, making conclusions for this parameter less definitive. Additionally, the exponents for $\theta$ and $D$ decline at the largest dataset sizes, deviating from the overall increasing trend. This likely results from a strong dependence of the test loss on the training scheme, suggesting potential for further network optimization. Therefore, more refined analysis may be needed to accurately characterize behavior in this large dataset regime. For comparison, the exponents for ResNet-18 and ViT, derived from the test loss data in Fig. 1, are also displayed in Fig. 5. These do not follow the same logarithmic pattern as inferred from FCN, despite involving the same task and their comparable model sizes.
@@
We observed that the average of MSE test loss, computed using either the arithmetic mean or median, exhibits scaling behaviors similar to the geometric mean: in both cases, the test loss decreases monotonically with increasing dataset size, with more pronounced reductions for larger models. Their scaling exponents show quantitative agreement with those from the geometric mean. This consistency can be attributed to the convergence of the geometric mean, arithmetic mean, and median for sufficiently large models [Figs. 3(b) and (d)]. Consequently, these measures can be used interchangeably for large models. However, adopting the geometric mean provides a more robust and consistent framework for evaluating model performance across all network sizes. Additionally, we have examined other activation functions, including ReLU, ELU, PReLU, and Swish. All results demonstrate power-law scaling behaviors consistent with the GeLu case. A key difference lies in the values of the scaling exponents, which deviate by at most 3% for $J$ and 28% for $D$. Further details are available in the Supplementary Information.
@@
In summary, we find that, for all network architectures, FCN, ResNet-18 and ViT, the test loss exhibits well-defined power-law scaling with dataset size $N_{D}$. Additionally, for FCN, increasing model parameters generally leads to higher scaling exponents.
@@
Figure 5: Power-law scaling exponents $\alpha_{D}$ as a function of FCN model size. Each marker displays the fitted parameter of $\alpha_{D}$ from the MSE test loss data in Fig. 4(a)–(c) for $\theta$, $J$, or $D$, respectively. The FCN model size $N_{M}$ ranges from approximately $8\times 10^{4}$ to $10^{7}$ parameters, corresponding to FCNs with $n_{n}$ from 4 to 512 and $n_{l}=3$. Dashed lines represent logarithmic fits of the form $\alpha_{D}=a_{D}\log_{10}(N_{M}/10^{6})+b_{D}$, where $a_{D}$ and $b_{D}$ are fitting parameters. The data points for the largest value of $N_{M}$ are omitted from the fit. Circle and diamond markers indicate the $\alpha_{D}$ values for ResNet-18 and ViT, respectively.